a horse runs along a circle with a speed of 20 km / h . a lantern is at the center of the circle . A fence is along the tangent to the circle at the point at which the horse starts . Find the speed with which the shadow of the horse moves along the fence at the moment when it covers 1/8 of the circle in km/h.

Grade:12

1 Answers

Arun
6 years ago
Dear student,
Draw a figure
 
tan @ = x/r
x = r* tan@
 
dx/dt = r sec²@ d@/dt = r*w* sec²@
= v sec ²@
 
Where @ = 2(pi)/8
So
dx/dt = 20 sec²(pi/4) = 20 *2 = 40 km/hr
 
Regards
Arun (askIITians forum expert)